OpenStudy (anonymous):

Write in sigma notation:4 + 6 + 8 + 10

OpenStudy (anonymous):

I know...nothing. So lost here...

OpenStudy (anonymous):

\[\sum_{n=2}^{n=5}(2n)\]

OpenStudy (anonymous):

Here : 4 + 6 + 8 + 10 So, it can be written as (they are multiples of 2) : \[2 \times 2, 3 \times 2, 4 \times 2, 5 \times 2\] If I let 2, 3, 4,5 as n, then: \[2 \times n, 2 \times n, 2 \times n, 2 \times n = 2n, \; 2n, \; 2n, \; 2n\] So here, \(n\) varies form 2 to 5: And Series is simply 2n: \[\huge \color{red}{\sum_{n = 2}^{n=5}(2n)}\]